# Modality Collapse as Mismatched Decoding: Information-Theoretic Limits of Multimodal LLMs

### v1

@@ 5.2 Modality-specific structure interferes with decoding
We directly test whether the decoder’s scoring rule functionally depends on modality-specific (MS) directions by projecting them out at inference time and measuring the change in cross-entropy. Unlike probing, which tests representational presence, this is a direct intervention on the decoder’s input.
We decompose the adapter output into eigenmodes (principal directions of variance) and ask: how much of this variance does the text distribution also use? Each eigenmode $u_{k}$ gets a text alignment score $\tilde{\alpha}(u_{k})=u_{k}^{\top}\Sigma_{T}u_{k}/\lambda_{k}$, the fraction of its variance that text representations share (§5.6). Modes with $\tilde{\alpha}<0.5$ are modality-specific (MS): most of their variance comes from the non-text modality. We then surgically remove selected directions from the decoder’s input ($z^{\prime}=z-\sum_{k\in S}(z^{\top}u_{k})\,u_{k}$) and measure the effect on cross-entropy. We compare three conditions (200 samples per model): (i) ablate all MS modes, (ii) ablate a matched number of top text-aligned (TA) modes, (iii) no ablation.
@@
Table 2: Causal ablation at the adapter output. Eigenmodes are classified as modality-specific (MS; $\tilde{\alpha}<0.5$) or text-aligned (TA), and matched subsets are projected out. Random controls ablate the same number of modes chosen uniformly at random (seed-averaged). For non-aligned encoders, removing MS modes improves loss while removing equal-sized TA subsets has negligible effect; the contrast is $160\times$ for Prismatic-D (t=25.2t=-25.2 vs. t=0.4t=-0.4). Random ablation gives intermediate effects proportional to variance removed. For SigLIP, few modes are MS and all effects are $1\%$.
MS structure is not merely unused; its presence degrades decoding. Unlike concept erasure methods such as LEACE (Belrose et al., 2023), which project out directions predictive of a labeled attribute, our ablation targets directions identified by distributional mismatch between modal and text covariances, without requiring any concept labels. The question is not whether a concept can be erased, but whether the decoder uses these distribution-mismatched directions at all.
Table 2 reveals a stark asymmetry across all three models. For Ultravox, removing 11 MS modes (63.6% of variance) slightly improves cross-entropy (-1.4%, t=7.0t=-7.0), while removing a matched TA subset has negligible effect (-0.3%, t=2.2t=-2.2). The Prismatic-D result is even more striking: ablating 53 MS modes (71% of variance) reduces loss by 11.1% (t=25.2t=-25.2), while ablating the top 47 TA modes (29% of variance) has no measurable effect (-0.07%, t=0.4t=-0.4, $p=0.69$). The decoder is functionally indifferent to the directions where most adapter variance concentrates, precisely as the mismatched-decoder theory predicts.
For Prismatic-S (text-aligned), only 14 of 100 modes are MS (24.6% of variance) and all ablation effects are below 1%, consistent with SigLIP producing representations the decoder already expects.
@@
The causal ablation establishes that MS directions interfere with decoding. We now isolate why some models have more MS variance than others, using the Prismatic pair as a controlled comparison that holds everything constant except encoder text alignment.
Both Prismatic VLMs share the same MLP adapter, Vicuna-7B backbone, and training recipe. The only difference is the vision encoder: DINOv2 (Oquab et al., 2024) (no text alignment) versus SigLIP (Zhai et al., 2023) (contrastive text–image loss). If our theory is correct, SigLIP should produce representations closer to $P_{\!T}$, yielding smaller $W_{\!1}(P_{\!M},P_{\!T})$ and less information loss.
The results confirm the prediction (Appendix Figure 1, Table 10). With DINOv2, visual information passes through the LLM essentially unchanged (-0.0% object category). With SigLIP, the same information types improve (++1.9% object category, ++2.3% super-category). This pattern replicates across modalities: in speech, Whisper-based encoders show dramatic lexical recovery (++92–95%) alongside speaker collapse (-8% to -39%); in LLaVA (CLIP encoder, text-aligned), all information types improve (Appendix Table 7).
5.4 The information accessibility gap in practice
The causal ablation (§5.2) establishes that the decoder does not use MS directions, and the controlled comparison (§5.3) traces this to encoder text alignment. Probes confirm the information has not been destroyed (Appendix Tables 75).
@@ 5.4 The information accessibility gap in practice
Three non-textual visual attributes (object count, size, spatial spread) derived from COCO annotations but absent from captions confirm the pattern in a milder mismatch regime ($L_{\log}\cdot W_{\!1}\approx 13$$54$ vs. $162$ for speech), explaining stagnation rather than degradation (Appendix Table 10).
@@ 5.5 Validating the bound
Theorem 2 predicts that GMI degradation scales with $L_{\log}\cdot W_{\!1}$. We estimate both quantities empirically (full results in Appendix Table 8). $L_{\log}$ is finite for all nine model-dataset combinations, confirming the Lipschitz log-score assumption (Assumption 2, Appendix A.2). Three patterns emerge: (i) $L_{\log}$ is content-dependent, decreasing as audio moves away from text ($9.08$ speech $\to$ $3.66$ environmental sounds in Ultravox); (ii) $L_{\log}$ varies $30\times$ across architectures ($0.29$ LLaVA to $9.08$ Ultravox); (iii) Prismatic-S has $2.8\times$ higher $L_{\log}$ than Prismatic-D ($1.12$ vs. $0.40$), consistent with the LLM being more responsive to text-aligned representations. $W_{\!1}$ ranges from 17.8 to 50.4 across models.
Gradient isotropy. The mean gradient norm along modality-specific directions ($g_{\text{MS}}$) is nearly identical to that along text-aligned directions ($g_{\text{TA}}$): $g_{\text{TA}}/g_{\text{MS}}=0.94\times$, Spearman $r=0.11$, $p=0.28$. This validates the scalar $L_{\log}$ assumption: the decoder responds with similar magnitude to perturbations in any direction. But MS modes carry structured variance that pushes predictions away from text distributions, while TA modes carry variance the decoder was trained to exploit. Same sensitivity, different semantic effect.
The product $L_{\log}\cdot W_{\!1}$ is largest for Ultravox ($162$), which shows the most dramatic degradation (-39% speaker), and smallest for LLaVA ($13.4$), where all information types improve. Prismatic-S has a higher product ($53.7$) than Prismatic-D ($20.2$) despite showing less degradation; this reflects higher decoder responsiveness to text-aligned representations rather than greater mismatch. The bound’s prefactor is also smaller for Prismatic-S because its effective support diameter is smaller ($D=16.2$ vs. $35.7$).
The raw bound uses the ambient representation diameter $D$, which makes the prefactor vacuously large. The support-restricted bound (Appendix A.3) allows replacing $D$ with the effective support diameter $D_{\text{eff}}$, estimated via the participation ratio of the pooled representation covariance. This yields $L_{\log}\cdot D_{\text{eff}}\in[3,7]$, making the prefactor moderate ($\sim$20–1100). The bound is structural rather than tight: it identifies $L_{\log}\cdot W_{\!1}$ as the governing quantity and correctly predicts which models collapse most.
Per-type prediction. The theory correctly predicts the aggregate pattern: text-aligned information recovers or improves, while non-text types degrade or stagnate. A finer prediction (degradation correlating with each type’s MS-variance share) does not hold at this granularity (Spearman $\rho=-0.40$, $p=0.60$), likely because the content-dependent $L_{\log}$ (which varies $2.5\times$ across audio types) modulates degradation beyond geometric alignment alone.
5.6 Mode alignment
The bound explains how much information is lost; mode alignment explains which directions carry the mismatch. We decompose the adapter output into its principal directions of variance (eigenmodes of $\Sigma_{M}$) and measure, for each direction, the fraction of its variance that the text distribution shares: $\tilde{\alpha}(u_{k})=u_{k}^{\top}\Sigma_{T}u_{k}/\lambda_{k}$. A score near 1 means the direction is shared with text; near 0 means it is unique to the modality (Appendix Table 11, Figure 2).
For non-aligned encoders, the largest direction (Mode 0) is modality-specific ($\tilde{\alpha}\leq 0.034$), capturing 51–79% of adapter variance. Most of what the adapter transmits is invisible to the text distribution. For SigLIP, even Mode 0 is shared with text ($\tilde{\alpha}=0.83$). As representations pass through the LLM, text-aligned variance is amplified (${\sim}150\times$) while MS variance remains flat: MS modes drop from 63.6% of adapter variance to <<1% at the final layer in Ultravox, not destroyed but drowned out.
@@ 6.1 Implications
The fix is objective-side, not encoder-side. If modality collapse is a consequence of the scoring rule being text-shaped, the natural intervention is to reshape the scoring rule. Low-rank adaptation (LoRA; Hu et al., 2022) on the LLM adjusts $q_{\psi}$ to better match $P_{\!M}$, effectively reducing the functional mismatch. Preliminary evidence from concurrent work (Billa, 2026) is consistent: applying LoRA to the LLM backbone reduces text dominance by 23.9%, while adapter-only training increases text dominance by 26.5%. This is predicted by Theorem 2: LoRA reduces the functional mismatch between $P_{\!M}$ and $P_{\!T}$ by reshaping the scoring rule, whereas adapter-only training leaves the decoder’s text-shaped sensitivity unchanged.
However, LoRA on the decoder is necessary but not sufficient: the training objective must explicitly target non-text information. We demonstrate this with two contrasting experiments on Ultravox, using the same LoRA configuration (r=16r{=}16, $\alpha{=}32$, targeting q/k/v/o_proj).
Negative example (text-centric objective). We apply a LoRA checkpoint trained for audio-text conflict resolution (ALME; Billa, 2026) and measure linear probe accuracy on LibriSpeech. Speaker identity accuracy is unchanged (55.0% vs. 54.5% base, within noise), as is lexical accuracy (53.5% vs. 52.6%). The ALME objective rewards resolving text-audio conflicts, not preserving speaker or acoustic attributes, so its gradient signal never reaches modality-specific directions.
Positive example (emotion objective). We train a LoRA on the LLM backbone with a forced-choice emotion detection objective on CREMA-D (6 emotions, 7,442 clips, standard causal LM loss on assistant tokens only). Generation accuracy on held-out samples improves from 17.3% to 61.8% (Table 3). Critically, the emotion probe at llm-final improves from .557.557 to .632.632 (++7.5%), while the speaker probe remains at chance (.135.135 vs. .137.137) and lexical accuracy remains near-ceiling (.994.994). Upstream layers (encoder, adapter) are identical across conditions, confirming the change is entirely in the LLM. The LoRA reshapes the scoring rule so that, at inference, the decoder responds to emotion-relevant directions while leaving others untouched, as Theorem 2 predicts: the emotion objective reduces $W_{\!1}$ along emotion-correlated dimensions without affecting orthogonal ones.
Table 3: LoRA intervention on Ultravox (CREMA-D, llm-final probe accuracy). The emotion LoRA selectively improves emotion accessibility without meaningfully affecting speaker or lexical probes. Generation accuracy measures forced-choice emotion detection on 1,002 held-out samples. The encoder and adapter outputs are identical across conditions (LoRA modifies only the LLM).
Condition Emotion Speaker Lexical Gen. acc.
Base Ultravox .557 .137 1.000 17.3%
@@
$\Delta$ ++7.5% -0.2% -0.6% ++44.5%
Together, the negative and positive examples confirm that the training objective determines which dimensions of the scoring rule are reshaped. A text-centric objective (ALME) leaves the decoder blind to non-text directions; an objective targeting a specific non-text attribute (emotion) teaches the decoder to respond to exactly that dimension.
Text-aligned encoders are a workaround, not a solution. Contrastive encoders like CLIP and SigLIP achieve low mismatch by mapping inputs onto text-correlated dimensions, but this is a projection, not a preservation: modality-specific information with no textual correlate is discarded at the encoder. The Prismatic comparison confirms this: SigLIP’s visual information “improves” through the LLM only because it was pre-projected onto text-like directions. CLIP gives the LLM access to textual correlates of visual information, not visual information itself.
Models must be explicitly trained to use non-text modalities. A text-trained decoder responds only to directions it was trained on (§5.6). No adapter can make it sensitive to directions it has never seen. To build models that genuinely leverage modality-specific information, not just its textual shadow, the training objective must reward extracting non-text directions, whether through LoRA, full multimodal pre-training, or other objective-side interventions. Without this, non-textual information remains hidden in plain sight.
The framework is architecture-agnostic. Our experiments use models with explicit adapter modules, but the analysis does not depend on them. The bound is a property of the decoder’s scoring rule, not of any particular architectural component. Any system where a text-trained decoder processes non-text representations faces the same constraint, whether the representations arrive through a linear projection, an MLP, a Q-Former, a discrete codebook (Borsos et al., 2023; Défossez et al., 2024), or no explicit adapter at all. The LLM’s own layers implicitly serve as the adapter. Discrete-token approaches add a quantization bottleneck that can only increase $W_{\!1}(P_{\!M},P_{\!T})$, making our bound conservative. The core mechanism is the mismatch between the decoder’s training distribution and its inputs at inference, and that mechanism is universal.
@@ 6.2 Limitations
Probes as information measures. Linear probes are necessary but not sufficient (Belinkov, 2022), but the key asymmetry (more information is linearly recoverable than the decoder uses) holds regardless. We estimate $L_{\log}$ via p95 gradient norms, not the true supremum. Additional limitations (scope, Assumption 1) are in Appendix D.
@@ 7 Conclusion
[…]
Building multimodal models that truly use what they perceive requires reshaping the decoder’s scoring rule. Text-aligned encoders appear to help but work by discarding non-textual information at the encoder. Our LoRA experiment shows the real fix: training with an emotion objective teaches the decoder to respond to emotion-relevant directions (++7.5% probe, ++44.5% generation accuracy) while leaving others untouched.
The framework extends beyond the models and modalities studied here. The bound is a property of the decoder’s scoring rule, not of any architectural component; no explicit adapter is required. At inference, every trained model has a fixed scoring rule. If that scoring rule was shaped predominantly by text, the constraint applies regardless of architecture or training recipe. The question is not whether modality-specific information can be encoded (it already is) but whether the training objective teaches the decoder to use it. Until it does, that information remains hidden in plain sight.
@@ Appendix A Theoretical framework: detailed treatment
This appendix provides the full formal development summarized in §4. We aim to make this self-contained: each result is preceded by an intuitive explanation, followed by the formal statement and a detailed proof with annotations.
@@ A.1 Generalized Mutual Information (GMI)
The core question. A representation $Z$ may contain information about a target $Y$ (e.g., the next token). But how much of that information can a specific decoder actually use? Standard mutual information $I(Z;Y)$ answers this for an ideal decoder that is perfectly tuned to the data it receives. A text-trained LLM is not ideal for non-text inputs: it was tuned on text representations, not speech or image representations. We need a measure that accounts for this mismatch. That measure is the Generalized Mutual Information (GMI) (Merhav et al., 1994; Scarlett et al., 2020).
We use the same notation as §3: $q_{\psi}(y|z,c)$ is the decoder’s next-token distribution (Eq. 1), $Z$ is the representation consumed by the LLM, $C$ is the text context, and $Y$ is the target token. Let $\ell_{\psi}(c,z,y):=\log q_{\psi}(y|z,c)$ be the decoder’s log-score (the log-probability it assigns to token $y$ given representation $z$ and context $c$), clipped at a floor $\eta=1/|\mathcal{V}|$ (where $\mathcal{V}$ is the decoder’s output vocabulary) to ensure boundedness.
Intuition. The GMI measures how well the decoder can tell the “right” representation from random alternatives. In words:
@@
When the decoder reliably scores the correct representation higher than random draws, the GMI is large. When it treats all representations interchangeably (i.e., it cannot tell which $Z$ goes with which $Y$), the GMI collapses to zero. This is closely related to the Donsker-Varadhan variational representation of KL divergence (Donsker & Varadhan, 1983), applied to the decoder’s scoring function rather than an arbitrary test function.
Formally, for a joint distribution $P$ over $(C,Z,Y)$ (either the text distribution $P_{\!T}$ or the modal distribution $P_{\!M}$ from §3), the GMI is:
$\mathrm{GMI}_{P}(q_{\psi})=\mathbb{E}_{(C,Z,Y)\sim P}\!\left[\ell_{\psi}(C,Z,Y)-\log\mathbb{E}_{Z^{\prime}\sim P(\cdot|C,Y)}\!\left[e^{\ell_{\psi}(C,Z^{\prime},Y)}\right]\right].$ (3)
@@
Under i.i.d. sampling from $P_{\!M}$ and standard measurability conditions, the maximum rate extractable by the fixed decoder $q_{\psi}$ is $R_{\mathrm{acc}}(P_{\!M},q_{\psi})=\mathrm{GMI}_{P_{\!M}}(q_{\psi})$.
This is a direct application of the random-coding argument of Scarlett et al. (2020), Chapters 2–3, with constant-composition codebooks under the GMI decoder metric. The key insight is that the decoder $q_{\psi}$ defines a fixed scoring rule (fixed at inference, regardless of how it was trained), and the maximum rate extractable under this rule, averaged over random codebooks drawn from $P_{\!M}$, converges to $\mathrm{GMI}_{P_{\!M}}(q_{\psi})$.
The implication is sharp: regardless of how much information a probe can extract from $Z$, the decoder can never exceed $\mathrm{GMI}_{P_{\!M}}(q_{\psi})$.

### v2

@@ 5.4 Modality-specific structure interferes with decoding
Probes show that non-text information is present; causal ablation tests whether the decoder is impacted by it. Unlike concept erasure methods such as LEACE (Belrose et al., 2023), which eliminate directions predictive of a labeled attribute, we ablate directions identified by distributional mismatch without concept labels. For each mode classified as MS ($\rho_{k}<0.5$) in §5.3, we remove its contribution from the representation: $z^{\prime}=z-\sum_{k\in S}(z^{\top}u_{k}),u_{k}$, and compare three conditions (200 samples per model): (i) ablate all MS modes, (ii) ablate a matched number of top TA modes, (iii) no ablation.
@@
Table 2: Causal ablation at the adapter output. Eigenmodes are classified as modality-specific (MS; $\rho_{k}<0.5$) or text-aligned (TA), and matched subsets are projected out. Random controls ablate the same number of modes chosen uniformly at random (seed-averaged). For non-aligned encoders, removing MS modes improves loss while removing equal-sized TA subsets has negligible effect; the contrast is $160\times$ for Prismatic-D (t=25.2t=-25.2 vs. t=0.4t=-0.4). Random ablation gives intermediate effects proportional to variance removed. For SigLIP, few modes are MS and all effects are <1{<}1%.
[…]
For Prismatic-S (text-aligned), only 14 of 100 modes are MS (24.6% of variance) and all ablation effects are below 1%, consistent with SigLIP representations already matching what the decoder expects.
@@
The causal ablation shows that MS directions influence decoding. Now, we explain why some models have more MS variability than others. The Prismatic pair is ideal for this, as it controls for all factors except the text-alignment of the encoder.
Both Prismatic VLMs use the same MLP adapter, Vicuna-7B backbone architecture and training protocol. The difference lies in the vision encoder: DINOv2 (Oquab et al., 2024) (no text alignment) versus SigLIP (Zhai et al., 2023) (contrastive text–image loss). According to our hypothesis, the representations learned by SigLIP should be more similar to $P_{\!T}$, leading to smaller $W_{\!1}(P_{\!M},P_{\!T})$ and less information loss.
This is what we see (Appendix Table 10): the split is sharper for non-textual attributes (object count, object size, spatial spread), which cannot be inferred from captions alone. For these features, SigLIP shows consistent improvement through the LLM (+4.5–8.2%) while DINOv2 gains less (+1.3–4.7%), with the largest gap on object count (+8.2% vs. +3.4%).
@@ 5.6 Validating the bound
$L_{\log}$ is finite for all nine model-dataset pairs, validating the Lipschitz assumption (Appendix Table 8). It varies with content ($9.08$ for speech to $3.66$ for environmental sounds in Ultravox) and differs $30\times$ across architectures ($0.29$ LLaVA to $9.08$ Ultravox). Prismatic-S has $2.8\times$ larger $L_{\log}$ than Prismatic-D ($1.12$ vs. $0.40$): the decoder is more responsive to representations that resemble text.
Gradient isotropy. To check whether the decoder selectively ignores MS directions at the gradient level, we compute the norm of the decoder’s loss gradient when the input representation is perturbed along each eigenmode direction. Let $g_{k}=|\nabla_{u_{k}}\mathcal{L}|$ be the gradient norm for mode $k$. The mean gradient norm for TA modes ($\bar{g}_{\text{TA}}$) and MS modes ($\bar{g}_{\text{MS}}$) are nearly equal ($\bar{g}_{\text{TA}}/\bar{g}_{\text{MS}}=0.94$, Spearman $r=0.11$, $p=0.28$): the decoder does not selectively ignore MS directions. This confirms the scalar $L_{\log}$ assumption.
The product $L_{\log}\cdot W_{\!1}$ ranks models by collapse severity. Ultravox ($L_{\log}\cdot W_{\!1}=162$) shows the most degradation (-39% speaker); LLaVA ($13.4$) shows none. Prismatic-S has a higher product ($53.7$) than Prismatic-D ($20.2$) despite less degradation, reflecting greater decoder responsiveness to text-aligned representations rather than greater mismatch. As noted in §4, the bound is an ordinal diagnostic: it correctly ranks models and predicts which information types degrade, even in the high-mismatch regime where it is numerically loose (Appendix A.3A.4).
@@ 6.1 Implications
The fix is objective-side, not encoder-side. If modality collapse is a result of the scoring rule being text-shaped, a logical solution would be to reshape the scoring rule. Low-rank adaptation (LoRA; Hu et al., 2022) on the LLM modifies $q$ to make it closer to $P_{\!M}$, effectively reducing the functional mismatch. Preliminary evidence from concurrent work on audio-text conflict supports this: applying LoRA to the LLM backbone decreases text dominance, whereas adapter-only training increases it. [^1] Citation omitted for double-blind review. This is predicted by Theorem 2: LoRA reduces the functional mismatch between $P_{\!M}$ and $P_{\!T}$ by reshaping the scoring rule, whereas adapter-only training does not change the decoder’s text-shaped sensitivity.
However, LoRA on the decoder is necessary but not sufficient: the training objective must explicitly target non-text information. We show this with a LoRA experiment on Ultravox (r=16r{=}16, $\alpha{=}32$, targeting q/k/v/o_proj).
LoRA with emotion objective. We fine-tune with LoRA on the LLM backbone using a forced-choice emotion detection objective on CREMA-D (6 emotions, 7,442 clips, standard causal LM loss on assistant tokens only). We evaluate the effect of LoRA on two measures: linear probe accuracy at the LLM’s final layer (measuring what information is accessible in the representation) and task accuracy (measuring whether the decoder actually uses it in its output). Task accuracy on held-out samples improves from 17.3% to 61.8% (Table 3). The emotion probe at llm-final improves from $557\pm.006$ to $632\pm.011$ (++7.5%), while the speaker probe remains at chance ($135\pm.006$ vs. $137\pm.008$) and lexical accuracy stays near-ceiling ($994\pm.001$). Upstream layers (encoder, adapter) are shared between conditions, so this is purely a change to the LLM. LoRA modifies the scoring rule such that at inference, the decoder will respond to the subset of directions that are emotion-relevant and leave the rest unchanged. Consistent with Theorem 2, the emotion objective reduces $W_{\!1}$ along the subset of directions correlated with emotion, while leaving all others unchanged.
Table 3: LoRA intervention on Ultravox (CREMA-D, llm-final probe accuracy). The emotion LoRA selectively improves emotion accessibility without meaningfully affecting speaker or lexical probes. Task accuracy measures forced-choice emotion detection on 1,002 held-out samples. The encoder and adapter outputs are identical across conditions (LoRA modifies only the LLM).
Condition Emotion (probe) Speaker (probe) Lexical (probe) Task acc.
Base Ultravox .557 .137 .999 17.3%
@@
$\Delta$ ++7.5% -0.2% -0.5% ++44.5%
Text-aligned encoders are a workaround, not a solution. Contrastive encoders like CLIP and SigLIP achieve low mismatch because they project the input data into the space of text-correlated features. The alignment is effective but selective in that it preserves what co-occurs with text and discards the rest. The Prismatic comparison shows that SigLIP’s visual representations “enhanced” by the LLM in fact only looked better because they were pre-projected into text-like directions. CLIP provides the LLM with visual features organized along text-correlated dimensions; genuinely visual information (texture, spatial layout) is discarded upstream.
Models must be explicitly trained to use non-text modalities. A text-only decoder will not listen to signals it was not trained on (§5.3). No adapter can make it sensitive to directions it has never seen. If we want a model to use modality-specific knowledge, rather than merely its text representation, we need to provide a training signal for it to do so, whether through LoRA, full multimodal pretraining, or some other objective-side intervention such as a modality-specific loss. Otherwise the information is just not available.
The framework is architecture-agnostic. The bound we present depends on the decoder’s scoring rule not the model’s architecture. The source of representations, a linear projection, an MLP, a Q-Former, a discrete codebook (Borsos et al., 2023; Défossez et al., 2024), or no explicit adapter at all (i.e., the LLM’s own layers act as the adapter), is immaterial. Discrete token methods introduce an additional quantization bottleneck that can only increase $W_{\!1}(P_{\!M},P_{\!T})$, making our bound more conservative. The essential phenomenon is the distributional shift the decoder sees between training and inference.
@@ 6.2 Limitations
Probes as information measures. Linear probes provide a necessary but not sufficient condition for information presence (Belinkov, 2022), but the main asymmetry (more information can be linearly recovered than the decoder uses) holds either way. We can only approximate $L_{\log}$ via p95 gradient norms, not compute the true supremum. Other caveats (in scope, Assumption 1) can be found in Appendix C.
@@
Scale. All five models studied are in the 7–8B parameter range. The bound (Theorem 2) depends on the scoring rule’s sensitivity ($L_{\log}$) and the distributional distance ($W_{\!1}$), not on parameter count, so the theoretical constraint is scale-invariant. For larger models the bound should remain consistent with our results insofar as a text-majority-training-objective paradigm remains. Validating the bound at the 70B scale is a direction for future work.
@@ 7 Conclusion
Modality collapse is caused by a failure of decoding, not encoding, as our experiments demonstrate. The GMI-Wasserstein bound shows theoretically that only information in the text-aligned directions is accessible to the decoder, and the rate at which information is lost depends on distributional divergence and decoder sensitivity.
To train multimodal models that properly exploit the information they receive, it is necessary to adjust the scoring function of the decoder. Text-aligned encoders do appear to improve performance, but they do so by ignoring all non-text modalities at the level of the encoder. Our LoRA experiment confirms this: an emotion-based objective trains the decoder to be sensitive to emotion-relevant messages (++7.5% probe, ++44.5% task accuracy), without affecting irrelevant messages. The question, for any multimodal model, is not whether modality-specific details can be represented (they are) but whether the training objective instructs the decoder to rely on them.
@@
Ethics & AI disclosure
This work analyses existing publicly available models and datasets; no new data was collected and no human subjects were involved. The author used Claude/Claude Code during preparation for manuscript critique, narrative feedback, literature search, and experiment implementation and debugging. All research design, theoretical development, experimental execution, analysis, and writing are the author’s own. The author takes full responsibility for all content.
@@ Appendix A Theoretical framework: detailed treatment
This appendix contains the detailed formal development outlined in §4. Each result is preceded by an informal intuitive statement followed by a complete proof.
@@ A.1 Generalized Mutual Information (GMI)
Overview. A representation $Z$ might encode some knowledge about a target $Y$ (e.g., the next token), but how much of this knowledge can a given decoder use? Standard mutual information $I(Z;Y)$ answers this for a decoder that is optimally matched to the representations it gets. A text-trained LLM is not optimally matched to non-text representations. It was matched to text representations, not speech representations, or image representations. The appropriate measure is the Generalized Mutual Information (GMI) (Merhav et al., 1994; Scarlett et al., 2020).
We reuse the notation from §3: $q(y|z,c)$ is the decoder’s next-token distribution (Eq. 1), $Z$ is the representation consumed by the LLM, $C$ is the text context, and $Y$ is the target token. Let $\ell(c,z,y):=\log q(y|z,c)$ be the decoder’s log-score (the log-probability it assigns to token $y$ given representation $z$ and context $c$), clipped at a floor $\eta=1/|\mathcal{V}|$ (where $\mathcal{V}$ is the decoder’s output vocabulary) to ensure boundedness.
Intuition. The GMI tells us how well the decoder can distinguish the “correct” representation from random alternatives:
@@
If the decoder consistently assigns a higher score to the true representation than to a uniformly sampled one, then the GMI is large. If the decoder cannot distinguish which $Z$ generated which $Y$ (assigning the same score to both), then the GMI is zero. This is similar to the Donsker-Varadhan variational representation of KL divergence (Donsker and Varadhan, 1983), but applied to the decoder’s scoring function instead of an arbitrary test function.
Formally, for a joint distribution $P$ over $(C,Z,Y)$ (which is either the text distribution $P_{\!T}$ or the modal distribution $P_{\!M}$ from §3), the GMI is given by:
$\mathrm{GMI}_{P}(q)=\mathbb{E}_{(C,Z,Y)\sim P}\!\left[\ell(C,Z,Y)-\log\mathbb{E}_{Z^{\prime}\sim P(\cdot|C,Y)}\!\left[e^{\ell(C,Z^{\prime},Y)}\right]\right].$ (3)
@@
Under i.i.d. sampling from $P_{\!M}$ and standard measurability conditions, the maximum rate extractable by the fixed decoder $q$ is $R_{\mathrm{acc}}(P_{\!M},q)=\mathrm{GMI}_{P_{\!M}}(q)$.
This is a classical result from the mismatched decoding literature; we state it here for completeness and to establish notation. The proof follows directly from the random-coding argument of Scarlett et al. (2020), Chapters 2–3, with constant-composition codebooks and the GMI decoder metric. The main idea is that the decoder $q$ defines a fixed scoring rule (fixed at inference, regardless of how it was trained), and the maximum rate extractable under this rule, averaged over random codebooks distributed according to $P_{\!M}$, approaches $\mathrm{GMI}_{P_{\!M}}(q)$.